40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Mon Mar 4 16:46:31 2013, yuta, Update, SUS, oplev calibration for PRM PRM_PIT.pngPRM_YAW.png
    Reply  Tue Mar 5 17:09:30 2013, yuta, Update, SUS, oplev calibration for ITMY ITMY_PIT.pngITMY_YAW.png
       Reply  Tue Apr 2 15:58:40 2013, Jenne, Update, SUS, oplev calibration for ITMX, BS opl_itmx.pngopl_bs.png
          Reply  Tue Apr 2 16:13:10 2013, rana, Update, SUS, oplev calibration for ITMX, BS 
             Reply  Tue Apr 2 16:55:34 2013, Jenne, Update, SUS, oplev calibration online for ITMs, BS, PRM comparison_opl.pdf
                Reply  Tue Apr 2 18:19:30 2013, Jenne, Update, SUS, BS, PRM oplev servos improved PRM_servo_matches_BS.pdfprm_bs_optical_levers_comparison.pdf
                Reply  Tue Apr 30 22:40:57 2013, Jenne, Update, SUS, oplev XY-plots reflect new calibration 
                   Reply  Sun Sep 22 22:05:10 2013, rana, Update, SUS, oplev XY-plots reflect new calibration noise.png
                      Reply  Thu Oct 3 01:06:48 2013, rana, jenne, Update, SUS, oplev XY-plots reflect new calibration etmx_ol.png
                         Reply  Thu Oct 3 09:49:49 2013, Steve, Update, SUS, ETMY oplev laser is sick 
                      Reply  Tue Nov 5 19:55:12 2013, rana, Update, SUS, oplev XY-plots reflect new calibration 
Message ID: 8389     Entry time: Tue Apr 2 15:58:40 2013     In reply to: 8232     Reply to this: 8390
Author: Jenne 
Type: Update 
Category: SUS 
Subject: oplev calibration for ITMX, BS 

[Jenne, Gabriele]

Optical lever calibrations:

ITMX pit calibration = -9.07 cts/mrad

ITMX yaw calibration = -12.33 cts/mrad

ITMX plot:opl_itmx.png

BS pit calibration = -22.86 cts/mrad

BS yaw calibration = -24.14 cts/mrad

BS plot:opl_bs.png

 

Method:  Similar to Manasa and Yuta's method last month.  We mounted each oplev QPD on a micrometer translation stage, centered the beam using the steering mirror, then used tdsavg to get 10 second averages of the _INMON channel for various settings of the micrometer stage.  For BS, we had to take out the PRM oplev to make room for the translation stage.  All QPDs were remounted in their original positions, within less than 1mm.  Measured the out-of-vac distances with the laser disto-meter, and the invac distances from the optic to the window from the CAD drawing.

Copying from other elog entries,

elog 8232:

We calibrated oplev for ITMY. Calibration factor for C1:SUS-ITMY_OL(PIT|YAW)_IN1 are;
  OLPIT: 6.29 +/- 0.11 counts/mrad
 
OLYAW: 5.74 +/- 0.09 counts/mrad

 

elog 8221:

We calibrated oplev for PRM. Calibration factor for C1:SUS-PRM_OL(PIT|YAW)_IN1 are;
  OLPIT: 15.6 +/- 0.3 counts/mrad
  OLYAW: 17.8 +/- 0.3 counts/mrad

ELOG V3.1.3-